OpenStudy (anonymous):

If I add three odd numbers and got the answer of 50, explain why this is wrong?

OpenStudy (solomonzelman):

3 odd integers will have an odd sum, just like any odd number of odd integers have an odd sum.

OpenStudy (anonymous):

Fro example 3+3 = 6. But 3+3+3=9. and 1+1+2 but 1+1+1= 3